log2(5x(x-8))=log2A+log2x+log2f(x)
A=5
f(x)=?

Answer by Theo(13342)

A = 5
log2(A) = log2(5)
log2(A) + log2(x) = log2(5) + log2(x) = log2(5x)
subtract log2(5x) from both sides of the equation to get:
log2(5x*(x-8)) - log2(5x) = log2(f(x))
this is equivalent to:
log2((5x*(x-8)/(5x)) = log2(f(x))
simplify to get log2(x-8) = log2(f(x))
this is true if and only if x-8 = f(x)
that's your solution.

you are using the properties of logarithms that state:

log(a) + log(b) = log(a*b)

log(a) - log(b) = log(a/b)